# Post-detection inference for sequential changepoint localization

### v4

@@ 3.2 An extension to composite pre-change
Note that (3.10b) is analogous to (3.6), while (3.10c) requires that the stopping time $\tau$ under change from $F_{0}^{*}$ to $F_{1}^{*}$ is stochastically larger than $\tau$ under change from $F_{0}$ to $F_{1}$, for any $(F_{0},F_{1})\in\mathcal{P}_{0}\times\mathcal{P}_{1}.$ To justify that (3.10c) is sensible, we note that the LFD pair $(F_{0}^{*},F_{1}^{*})$ is the closest pair within $\mathcal{P}_{0}\times\mathcal{P}_{1}$ (according to various metrics or divergences such as Kullback-Leibler), i.e., it renders the change detection problem most difficult. Therefore, it is reasonable to assume that $\tau$ under change from $F_{0}^{*}$ to $F_{1}^{*}$ is stochastically largest. (3.10a) can be justified similarly because $F_{0}^{*}$ is the “closest” element of the pre-change class to the post-change class.
Theorem 3.9.
@@
In the setting described above, for any change detection method with stop time $\tau$ that is well-behaved in the sense of (3.10a) and $p_{T}\neq 0$,
@@
$\limsup_{N\to\infty}\sup_{\begin{subarray}{c}F_{0}\in\mathcal{P}_{0}\\ F_{1}\in\mathcal{P}_{1}\end{subarray}}\mathbb{E}_{F_{0},T,F_{1}}(|{\mathcal{C}}|\mid\tau\geq T)\leq\frac{2^{s_{0}}}{\alpha^{s_{0}}p_{{T}}^{s_{0}+1}}\frac{1-\rho_{0}(s_{0})^{T-1}}{1-\rho_{0}(s_{0})}+1+\Delta_{\mathcal{P}_{0},T,\mathcal{P}_{1}},$ (3.11)
@@
where $\Delta_{\mathcal{P}_{0},T,\mathcal{P}_{1}}:=\displaystyle\sup_{\begin{subarray}{c}F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}\end{subarray}}\Delta_{F_{0},T,F_{1}}$. Additionally, if the stopping time $\tau$ is well-behaved in the sense of (3.10b) and (3.10c) as well, the last term in (3.11) can be improved to $\min\left\{\frac{1}{p_{T}}\left(\left(\frac{2}{\alpha}\right)^{s_{1}}\times\frac{\rho_{1}(s_{1})^{s_{1}}}{1-\rho_{1}(s_{1})^{s_{1}}}+\frac{\rho_{1}(s_{1})}{1-\rho_{1}(s_{1})}\right),\Delta_{\mathcal{P}_{0},T,\mathcal{P}_{1}}\right\}.$
@@ 4.1 Known pre- and post-change distributions
We develop a simulation-based scheme to approximate the exact $\alpha$-th quantile of the distribution of the test statistic $M_{t}$. Here, $M_{t}$ need not be of the form (3.4), it can be any test statistic based on $X_{1},\cdots,X_{\tau}$. Assuming known pre- and post-change distributions, we draw $B$ many independent streams of data under $H_{0,t}$. For the $j$-th sequence, denoted as $\{X_{n}^{j}\}_{n}$, we simulate data points until a change is detected at $\tau_{j}^{\prime}=\mathcal{A}(\{X_{n}^{j}\}_{n})$ or until a fixed time $L$, whichever happens first. Then, we compute a test statistic $M_{t,L}^{j}$, based on $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$ (details are described in Supplementary Section B). Our confidence set is the collection of all $t\in\{1,\cdots,\tau\}$, for which we fail to reject $H_{0,t}$:
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-\alpha r_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})\right\},$ (4.1)
@@
8:  Compute test statistic $M_{t}$ (e.g. (3.2)) based on $\tau$ and $X_{1},\dots,X_{\tau}$
@@
Theorem 4.1.
For any fixed $\alpha\in(0,1)$, $B,N,T\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$, and change detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}_{1}$ defined in (4.1), constructed using any data-independent unbiased estimator $r_{t}$ of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-{\alpha}.$
Remark 4.2.
@@ 4.2 Adaptive confidence set for known pre-change distribution and parametric post-change class
Since the post-change parameter is unknown, for each $t$, we estimate it based on $X_{t},\cdots,X_{\tau}$. However, the sample size used for the estimate might be too small to have a reliable estimate. To address this issue, we would use interval estimates instead of point estimates, allowing us to achieve a predetermined control over the error. Notably, conventional confidence intervals are invalid here, since $\tau$ is random. Therefore, we rely on confidence sequences [4, 12]. A confidence sequence having coverage 1c1-c is defined as a sequence of confidence sets $\{\operatorname{CS}(X_{1},\cdots,X_{n};1-c)\}_{n\in\mathbb{N}}$, satisfying the uniform guarantee:
$\mathbb{P}(\forall n\in\mathbb{N},\theta\in\operatorname{CS}(X_{1},\cdots,X_{n};1-c))\geq 1-c.$ (4.2)
@@
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}\left(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}\right)\right\}.$ (4.3)
Algorithm 2 summarizes the method. Note that, even when $\mathcal{S}^{\prime}_{\tau-t+1}$ is an infinite set, the threshold $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ can be approximated by discretizing $\mathcal{S}^{\prime}_{\tau-t+1}$ into a grid and evaluating the quantile for values of $\theta$ on this grid, taking their maximum. While this approach is conceptually straightforward, it is computationally intensive; related experiments are in Section F of the Supplementary. Moreover, in many cases, we can derive a suitable upper bound of this threshold, which is computationally, statistically efficient and does not require any approximation. Details are provided in Section C of the Supplementary, along with two concrete examples of the Gaussian and Poisson mean change problems.
@@
Theorem 4.3.
Let $\mathcal{C}_{1}$ be the confidence set for changepoint $T\in\mathbb{N}$ as defined in (4.3), constructed using any data-independent unbiased estimator $r_{t}$ of $\mathbb{P}_{\theta_{0},\infty}[\tau\geq t]$, for some $\alpha,\beta\in(0,1)$, $N,B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$. Then for any given detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$ and any $\theta_{1}\in\Theta_{1}$, $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta.$
@@ 4.3 Adaptive confidence set for parametric pre- and post-change classes
Since both pre- and post-change parameters are unknown, for each $t$, we estimate them based on $X_{1},\cdots,X_{t-1}$ and $X_{t},\cdots,X_{\tau}$ respectively. Instead of point estimates, we rely on interval estimates to achieve a predetermined control over the error. Suppose that we have some confidence interval (CI) procedure $\operatorname{CI}(X_{1},\cdots,X_{n};1-c)$ that constructs CI using $\{X_{i}\}_{i=1}^{n}$ for $\theta_{0}$ with coverage 1c1-c and a confidence sequence (CS) procedure $\{\operatorname{CS}(X_{1},\cdots,X_{n};1-c)\}_{n}$, that constructs CS using $\{X_{i}\}_{i=1}^{n}$ for $\theta_{1}$ having coverage 1c1-c. We construct $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t}^{*})$ and $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t}^{*})$, where $r_{t}^{*}$ is computed under Assumption 3.5.
For each $j=1,\cdots,B$, for some fixed $B\in\mathbb{N}$, suppose we have sequences $\{X^{j}_{n}(\theta,t,\theta^{\prime})\}_{n}$ having joint distribution as $\mathbb{P}_{\theta,t,\theta^{\prime}}$. Note that the sequences must be independent as $j$ varies. However, for any fixed $j$, the i.i.d. sequences $\{X^{j}_{n}(\theta,t,\theta^{\prime})\}_{n}$ can be dependent across $t$ and $\theta\in\mathcal{S}_{t-1},\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$. A specific method for generating these sequences is detailed in Section D of the Supplementary. We draw the data until a change is detected at the stopping time $\tau^{j,t}_{\theta,\theta^{\prime}}$ using $\mathcal{A}$ or until time $L$, whichever occurs first. Then, we compute the same test statistics truncated at $L$, based on $\tau^{j,t}_{\theta,\theta^{\prime}}$ and $X^{j}_{1}(\theta,t,\theta^{\prime}),\cdots,X^{j}_{\tau^{j,t}_{\theta,\theta^{\prime}}\wedge L}(\theta,t,\theta^{\prime})$, which is denoted as $M^{j}_{t,L}(\theta,\theta^{\prime})$. Similar to (4.3), the adaptive confidence set is defined as
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{\mathbf{S}}^{(t)}}\operatorname{Quantile}(1-{\alpha}r_{t}^{*};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})\right\}$ (4.4)
where $\mathcal{\mathbf{S}}^{(t)}=\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$. Algorithm 3 contains an overview. If the detection algorithm controls PFA, a simpler algorithm can be used, which avoids Assumption 3.5 and provides coverage guarantees, as elaborated in Section G of the Supplementary.
As before, one can always adopt a grid-based approach, which is conceptually straightforward, but it is computationally intensive; corresponding experimental results are reported in Section F of the Supplementary. Additionally, in many cases, we can derive a suitable upper bound of this threshold, which is computationally efficient and exact, but that can be a bit more conservative. We provide the details in Section D of the Supplementary, along with two concrete examples.
@@
Theorem 4.4.
Let $\mathcal{C}_{1}$ be the confidence sets for changepoint $T\in\mathbb{N}$ defined in (4.4), constructed using any data-independent unbiased estimator $r_{t}^{*}$ of $\mathbb{P}_{\theta_{0}^{*},\infty}[\tau\geq t]$, for some $\alpha,\beta,\gamma\in(0,1)$, $N,B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$. For any $\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1}$, and any change detection algorithm $\mathcal{A}$ satisfying Assumption 3.5 and having $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$, the following guarantee holds: $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta-\gamma.$
Remark 4.5.
As a special case, when pre- and post-change parameters are known, i.e., $\Theta_{i}=\{\theta_{i}\}$ are singleton sets, $i=0,1$, we have $\mathcal{S}=\{\theta_{0}\},\mathcal{S}^{\prime}=\{\theta_{1}\}$ with $\beta=0$. So, in this case, Algorithms 3 and 2 reduces to our algorithm designed for the known pre- and post-change setting, and similarly, Theorems 4.3 and 4.4 reduces to Theorem 4.1.

### v6

@@ 3.2 An extension to composite pre-change
Note that (15b) is analogous to (11), while (15c) requires that the stopping time $\tau$ under change from $F_{0}^{*}$ to $F_{1}^{*}$ is stochastically larger than $\tau$ under change from $F_{0}$ to $F_{1}$, for any $(F_{0},F_{1})\in\mathcal{P}_{0}\times\mathcal{P}_{1}.$ To justify that (15c) is sensible, we note that the LFD pair $(F_{0}^{*},F_{1}^{*})$ is the closest pair within $\mathcal{P}_{0}\times\mathcal{P}_{1}$ (according to various metrics or divergences such as Kullback-Leibler), i.e., it renders the change detection problem most difficult. Therefore, it is reasonable to assume that $\tau$ under change from $F_{0}^{*}$ to $F_{1}^{*}$ is stochastically largest. (15a) can be justified similarly because $F_{0}^{*}$ is the “closest” element of the pre-change class to the post-change class.
Theorem 5.
@@
In the setting described above, for any change detection method with stop time $\tau$ that is well-behaved in the sense of (15a) and $p_{T}\neq 0$,
@@
$\limsup_{N\to\infty}\sup_{\begin{subarray}{c}F_{0}\in\mathcal{P}_{0}\\ F_{1}\in\mathcal{P}_{1}\end{subarray}}\mathbb{E}_{F_{0},T,F_{1}}(|{\mathcal{C}}|\mid\tau\geq T)\leq\frac{2^{s_{0}}}{\alpha^{s_{0}}p_{{T}}^{s_{0}+1}}\frac{1-\rho_{0}(s_{0})^{T-1}}{1-\rho_{0}(s_{0})}+1+\Delta_{\mathcal{P}_{0},T,\mathcal{P}_{1}},$ (16)
@@
where $\Delta_{\mathcal{P}_{0},T,\mathcal{P}_{1}}:=\displaystyle\sup_{\begin{subarray}{c}F_{0}\in\mathcal{P}_{0},F_{1}\in\mathcal{P}_{1}\end{subarray}}\Delta_{F_{0},T,F_{1}}$. Additionally, if the stopping time $\tau$ is well-behaved in the sense of (15b) and (15c) as well, the last term in (16) can be improved to
$\min\left\{\frac{1}{p_{T}}\left(\left(\frac{2}{\alpha}\right)^{s_{1}}\times\frac{\rho_{1}(s_{1})^{s_{1}}}{1-\rho_{1}(s_{1})^{s_{1}}}+\frac{\rho_{1}(s_{1})}{1-\rho_{1}(s_{1})}\right),\Delta_{\mathcal{P}_{0},T,\mathcal{P}_{1}}\right\}.$
@@ 4.1 Known pre- and post-change distributions
We develop a simulation-based scheme to approximate the exact $\alpha$-th quantile of the distribution of the test statistic $M_{t}$. Here, $M_{t}$ need not be of the form (9), it can be any test statistic based on $X_{1},\cdots,X_{\tau}$. Assuming known pre- and post-change distributions, we draw $B$ many independent streams of data under $H_{0,t}$. For the $j$-th sequence, denoted as $\{X_{n}^{j}\}_{n}$, we simulate data points until a change is detected at $\tau_{j}^{\prime}=\mathcal{A}(\{X_{n}^{j}\}_{n})$ or until a fixed time $L$, whichever happens first. Then, we compute a test statistic $M_{t,L}^{j}$, based on $X_{1}^{j},\cdots,X_{\tau_{j}^{\prime}\wedge L}^{j}$ (details are described in Supplementary Section B). Our confidence set is the collection of all $t\in\{1,\cdots,\tau\}$, for which we fail to reject $H_{0,t}$:
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\operatorname{Quantile}(1-\alpha r_{t};M_{t},M_{t,L}^{1},\cdots,M_{t,L}^{B})\right\},$ (17)
@@
8:  Compute test statistic $M_{t}$ (e.g. (8)) based on $\tau$ and $X_{1},\dots,X_{\tau}$
@@
Theorem 6.
For any fixed $\alpha\in(0,1)$, $B,N,T\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$, and change detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{F_{0},\infty}(\tau\geq T)\neq 0$, the confidence set $\mathcal{C}_{1}$ defined in (17), constructed using any data-independent unbiased estimator $r_{t}$ of $\mathbb{P}_{F_{0},\infty}[\tau\geq t]$, satisfies $\mathbb{P}_{F_{0},T,F_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-{\alpha}.$
Remark 2.
@@ 4.2 Adaptive confidence set for known pre-change distribution and parametric post-change class
Since the post-change parameter is unknown, for each $t$, we estimate it based on $X_{t},\cdots,X_{\tau}$. However, the sample size used for the estimate might be too small to have a reliable estimate. To address this issue, we would use interval estimates instead of point estimates, allowing us to achieve a predetermined control over the error. Notably, conventional confidence intervals are invalid here, since $\tau$ is random. Therefore, we rely on confidence sequences [darling1967confidence, howard2021time]. A confidence sequence having coverage 1c1-c is defined as a sequence of confidence sets $\{\operatorname{CS}(X_{1},\cdots,X_{n};1-c)\}_{n\in\mathbb{N}}$, satisfying the uniform guarantee:
$\mathbb{P}(\forall n\in\mathbb{N},\theta\in\operatorname{CS}(X_{1},\cdots,X_{n};1-c))\geq 1-c.$ (18)
@@
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}\left(1-\alpha r_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B}\right)\right\}.$ (19)
Algorithm 2 summarizes the method. Note that, even when $\mathcal{S}^{\prime}_{\tau-t+1}$ is an infinite set, the threshold $\sup_{\theta\in\mathcal{S}^{\prime}_{\tau-t+1}}\operatorname{Quantile}(1-\hat{\alpha}_{t};M_{t},\{M_{t,L}^{j}(\theta)\}_{j=1}^{B})$ can be approximated by discretizing $\mathcal{S}^{\prime}_{\tau-t+1}$ into a grid and evaluating the quantile for values of $\theta$ on this grid, taking their maximum. While this approach is conceptually straightforward, it is computationally intensive. Moreover, in many cases, we can derive a suitable upper bound of this threshold, which is computationally, statistically efficient and does not require any approximation. Details are provided in Section C of the Supplementary, along with a concrete example of the Gaussian mean change problems.
@@
Theorem 7.
Let $\mathcal{C}_{1}$ be the confidence set for changepoint $T\in\mathbb{N}$ as defined in (19), constructed using any data-independent unbiased estimator $r_{t}$ of $\mathbb{P}_{\theta_{0},\infty}[\tau\geq t]$, for some $\alpha,\beta\in(0,1)$, $N,B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$. Then for any given detection algorithm $\mathcal{A}$ with $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$ and any $\theta_{1}\in\Theta_{1}$, $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta.$
@@ 4.3 Adaptive confidence set for parametric pre- and post-change classes
Since both pre- and post-change parameters are unknown, for each $t$, we estimate them based on $X_{1},\cdots,X_{t-1}$ and $X_{t},\cdots,X_{\tau}$ respectively. Instead of point estimates, we rely on interval estimates to achieve a predetermined control over the error. Suppose that we have some confidence interval (CI) procedure $\operatorname{CI}(X_{1},\cdots,X_{n};1-c)$ that constructs CI using $\{X_{i}\}_{i=1}^{n}$ for $\theta_{0}$ with coverage 1c1-c and a confidence sequence (CS) procedure $\{\operatorname{CS}(X_{1},\cdots,X_{n};1-c)\}_{n}$, that constructs CS using $\{X_{i}\}_{i=1}^{n}$ for $\theta_{1}$ having coverage 1c1-c.
We construct $\mathcal{S}_{t-1}=\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\gamma r_{t}^{*})$ and $\mathcal{S}^{\prime}_{\tau-t+1}=\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\beta r_{t}^{*})$, where $r_{t}^{*}$ is computed under Assumption 1.
For each $j=1,\cdots,B$, for some fixed $B\in\mathbb{N}$, suppose we have sequences $\{X^{j}_{n}(\theta,t,\theta^{\prime})\}_{n}$ having joint distribution as $\mathbb{P}_{\theta,t,\theta^{\prime}}$. Note that the sequences must be independent as $j$ varies. However, for any fixed $j$, the i.i.d. sequences $\{X^{j}_{n}(\theta,t,\theta^{\prime})\}_{n}$ can be dependent across $t$ and $\theta\in\mathcal{S}_{t-1},\theta\in\mathcal{S}^{\prime}_{\tau-t+1}$. A specific method for generating these sequences is detailed in Section D of the Supplementary. We draw the data until a change is detected at the stopping time $\tau^{j,t}_{\theta,\theta^{\prime}}$ using $\mathcal{A}$ or until time $L$, whichever occurs first. Then, we compute the same test statistics truncated at $L$, based on $\tau^{j,t}_{\theta,\theta^{\prime}}$ and $X^{j}_{1}(\theta,t,\theta^{\prime}),\cdots,X^{j}_{\tau^{j,t}_{\theta,\theta^{\prime}}\wedge L}(\theta,t,\theta^{\prime})$, which is denoted as $M^{j}_{t,L}(\theta,\theta^{\prime})$.
Similar to (19), the adaptive confidence set is defined as
$\mathcal{C}_{1}=\left\{t\in\mathbb{N}:t\leq\tau,M_{t}\leq\sup_{(\theta,\theta^{\prime})\in\mathcal{\mathbf{S}}^{(t)}}\operatorname{Quantile}(1-{\alpha}r_{t}^{*};M_{t},\{M_{t,L}^{j}(\theta,\theta^{\prime})\}_{j=1}^{B})\right\}$ (20)
where $\mathcal{\mathbf{S}}^{(t)}=\mathcal{S}_{t-1}\times\mathcal{S}^{\prime}_{\tau-t+1}$. Algorithm 3 contains an overview. If the detection algorithm controls PFA, a simpler algorithm can be used, which avoids Assumption 1 and provides coverage guarantees, as elaborated in Section F of the Supplementary.
As before, one can always adopt a grid-based approach, which is conceptually straightforward, but it is computationally intensive; corresponding experimental results are reported in Section F of the Supplementary. Additionally, in many cases, we can derive a suitable upper bound of this threshold, which is computationally efficient and exact, but that can be a bit more conservative. We provide the details in Section D of the Supplementary, along with a concrete example.
@@
Theorem 8.
Let $\mathcal{C}_{1}$ be the confidence sets for changepoint $T\in\mathbb{N}$ defined in (20), constructed using any data-independent unbiased estimator $r_{t}^{*}$ of $\mathbb{P}_{\theta_{0}^{*},\infty}[\tau\geq t]$, for some $\alpha,\beta,\gamma\in(0,1)$, $N,B\in\mathbb{N}$, $L\in\mathbb{N}\cup\{\infty\}$. For any $\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1}$, and any change detection algorithm $\mathcal{A}$ satisfying Assumption 1 and having $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$, the following guarantee holds: $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}_{1}\mid\tau\geq T)\geq 1-\alpha-\beta-\gamma.$
Remark 3.
As a special case, when pre- and post-change parameters are known, i.e., $\Theta_{i}=\{\theta_{i}\}$ are singleton sets, $i=0,1$, we have $\mathcal{S}=\{\theta_{0}\},\mathcal{S}^{\prime}=\{\theta_{1}\}$ with $\beta=0$. So, in this case, Algorithms 3 and 2 reduces to our algorithm designed for the known pre- and post-change setting, and similarly, Theorems 7 and 8 reduces to Theorem 6.